# MGMCpaper.bib

### v2

@@ A.6 Bounds on the IACT
To derive (4.28) note that if $\theta_{L}^{(0)}\sim\mathcal{N}(\mu_{L},A_{L}^{-1})$ then according to Theorem 4.1 $\theta_{L}^{(m)}\sim\mathcal{N}(\mu_{L},A_{L}^{-1})$ for all $m\geq 0$. With (4.26c) in Lemma 4.8 this implies that
@@
Inserting this into (4.27) gives the desired result. To derive a bound on $\tau_{\mathrm{int},F}$, i.e. to show the first inequality in (4.29), observe that the individual terms in the sum in (4.28) can be bound as follows if we set $w:=A_{L}^{-1/2}F_{L}$
$\frac{F_{L}^{\top}X_{L}^{s}A_{L}^{-1}F_{L}}{F_{L}^{\top}A_{L}^{-1}}$ $\leq\left|\frac{F_{L}^{\top}X_{L}^{s}A_{L}^{-1}F_{L}}{F_{L}^{\top}A_{L}^{-1}F_{L}^{\top}}\right|=\left|\frac{w^{\top}A_{L}^{1/2}X_{L}^{s}A_{L}^{-1/2}w}{w^{\top}w}\right|\leq||A_{L}^{1/2}X_{L}^{s}A_{L}^{-1/2}||_{2}\leq||X_{L}||_{A_{L}}^{s},$ (A.32)
where the energy norm $||\cdot||_{A_{L}}$ is defined in (4.19). Using (A.32) and summing the geometric series we get the desired bound on $\tau_{\mathrm{int},F}$:
$\tau_{\mathrm{int},F_{L}}$ $=1+2\sum_{s=1}^{\infty}\frac{F_{L}^{\top}X_{L}^{s}A_{L}^{-1}F_{L}}{F_{L}^{\top}A_{L}^{-1}F_{L}}\leq 1+2\sum_{s=1}^{\infty}||X_{L}||_{A_{L}}^{s}=\frac{1+||X_{L}||_{A_{L}}}{1-||X_{L}||_{A_{L}}}.$ (A.33)
@@
$\tau_{\mathrm{int},F_{L}}$ $=1+2\sum_{s=1}^{\infty}\frac{\alpha_{s}}{\beta},$ $\tau_{\mathrm{int},F_{L}^{\top}}^{(m)}$ $=1+2\sum_{s=1}^{\infty}\frac{\alpha^{\prime}_{m,s}}{\beta^{\prime}_{m}}=\tau_{\mathrm{int},F_{L}}+2\sum_{s=1}^{\infty}\left(\frac{\alpha^{\prime}_{m,s}}{\beta^{\prime}_{m}}-\frac{\alpha_{s}}{\beta}\right).$
@@
holds. Hence, with $w:=A_{L}^{-1/2}F_{L}$ and since $\left|\left|A_{L}^{1/2}X_{L}^{m}A_{L}^{-1/2}\right|\right|_{2}=\left|\left|A_{L}^{-1/2}(X_{L}^{\top})^{m}A_{L}^{1/2}\right|\right|_{2}=||X_{L}||^{m}_{A_{L}}$
$\left|\frac{\beta^{\prime}_{m}-\beta}{\beta}\right|$ $=\left|\frac{F_{L}^{\top}(C_{L}^{(m)}-A_{L}^{-1})F_{L}}{F_{L}^{\top}A_{L}^{-1}F_{L}}\right|=\left|\frac{w^{\top}A_{L}^{1/2}(C_{L}^{(m)}-A_{L}^{-1})A_{L}^{1/2}w}{w^{\top}w}\right|\leq\left|\left|A_{L}^{1/2}(C_{L}^{(m)}-A_{L}^{-1})A_{L}^{1/2}\right|\right|_{2}$ (A.38)
$=\left|\left|\left(A_{L}^{1/2}X_{L}^{m}A_{L}^{-1/2}\right)\left(A_{L}^{1/2}(C_{L}^{(0)}-A_{L}^{-1})A_{L}^{1/2}\right)\left(A_{L}^{-1/2}(X_{L}^{\top})^{m}A_{L}^{1/2}\right)\right|\right|_{2}$
$\leq\left|\left|I-A_{L}^{1/2}C_{L}^{(0)}A_{L}^{1/2}\right|\right|_{2}||X_{L}||_{A_{L}}^{2m}\leq(1+C_{0})||X_{L}||_{A_{L}}^{2m}$
@@
$\leq\left|\left|A_{L}^{1/2}X_{L}^{s}A_{L}^{-1/2}A_{L}^{1/2}(C_{L}^{(m)}-A_{L}^{-1})A_{L}^{-1/2}A_{L}\right|\right|_{2}$
$=\left|\left|\left(A_{L}^{1/2}X_{L}^{s}A_{L}^{-1/2}\right)\left(A_{L}^{1/2}X_{L}^{m}A_{L}^{-1/2}\right)\left(A_{L}^{1/2}(C_{L}^{(0)}-A_{L}^{-1})A_{L}^{1/2}\right)\left(A_{L}^{-1/2}(X_{L}^{\top})^{m}A_{L}^{1/2}\right)\right|\right|_{2}$
$\leq\left|\left|I-A_{L}^{1/2}C_{L}^{(0)}A_{L}^{1/2}\right|\right|_{2}||X_{L}||_{A_{L}}^{2m+s}\leq(1+C_{0})||X_{L}||_{A_{L}}^{2m+s}.$
Similar arguments show that $|\alpha_{s}/\beta|\leq||X_{L}||^{s}_{A_{L}}$.
@@
$\left|\frac{\alpha^{\prime}_{m,s}}{\beta^{\prime}_{m}}-\frac{\alpha_{s}}{\beta}\right|$ $\leq\frac{\beta}{\beta^{\prime}_{m}}\left(\left|\frac{\alpha^{\prime}_{m,s}-\alpha_{s}}{\beta}\right|+\left|\frac{\alpha_{s}}{\beta}\right|\cdot\left|\frac{\beta^{\prime}_{m}-\beta}{\beta}\right|\right)\leq 2(1+C_{0})\frac{||X_{L}||_{A_{L}}^{2m+s}}{1-||X_{L}||_{A_{L}}^{2m}}\qquad\text{for $m\geq 1$}.$ (A.41)
@@
$|\tau_{\mathrm{int},F_{L}}^{(m)}-\tau_{\mathrm{int},F_{L}}|$ $\leq 2\sum_{s=1}^{\infty}\left|\frac{\alpha^{\prime}_{m,s}}{\beta^{\prime}_{m}}-\frac{\alpha_{s}}{\beta}\right|\leq 4(1+C_{0})\frac{||X_{L}||_{A_{L}}^{2m}}{1-||X_{L}||_{A_{L}}^{2m}}\sum_{s=1}^{\infty}||X_{L}||_{A_{L}}^{s}$ (A.42)
$=C\frac{||X_{L}||_{A_{L}}^{2m+1}}{(1-||X_{L}||_{A_{L}})(1-||X_{L}||_{A_{L}}^{2m})}\qquad\text{for $m\geq 1$},$
with the grid-independent constant $C:=4(1+C_{0})$. We conclude that $\tau_{\mathrm{int},F_{L}}^{(m)}$ converges to $\tau_{\mathrm{int},F_{L}}$ for $m\rightarrow\infty$. ∎
@@ A.8 Approximation property for perturbed matrix
where we used Assumption 2.1.
@@
$a(z_{\ell},\tilde{u}_{\ell})+b(z_{\ell},\tilde{u}_{\ell})$ $=f_{\ell}^{\top}P_{\ell}^{-1}z_{\ell}=\langle(P_{\ell}^{-1})^{*}f_{\ell},z_{\ell}\rangle_{H}=F_{\ell}(z_{\ell})\qquad\text{ for all }z_{\ell}\in V_{\ell},$
@@
Finally, Assumption 2.1 implies $\|P_{\ell}^{-1}\varphi_{\ell}\|_{2}\leq\frac{c_{2}}{\Phi(\ell)}\|\varphi_{\ell}\|_{H}$ for $\varphi_{\ell}\in V_{\ell}$ and thus
@@ A.9 Optimality of MGMC
Finally, noting that $u:=\widetilde{A}^{-1}f$ satisfies the problem $a(u,v)+b(u,v)=\langle f,v\rangle_{H}$ for all $v\in H$, Assumption 2.2 implies that $P_{L}u_{L}:=P_{L}\widetilde{A}_{L}^{-1}f_{L}$ with $f_{L}=(\langle f,\phi_{j}^{L}\rangle_{H})_{j=1,\dots,n_{L}}$ converges to $u$ in $V$ as $L\to\infty$. Hence,
@@
Similarly, $\lim_{L\to\infty}\mathrm{Cov}(F_{L}^{\top}\theta_{L})=\lim_{L\to\infty}\mathcal{F}P_{L}\widetilde{A}_{L}^{-1}F_{L}=\lim_{L\to\infty}\mathcal{F}\widetilde{A}^{-1}\mathcal{F}$, where we used $F_{L}=(\langle\mathcal{F},\phi_{j}^{L}\rangle_{H})_{j=1,\dots,n_{L}}$. Thus, the characteristic function of $(F_{L}^{\top}\theta_{L}^{*})_{L}$ converges to that of $(\widetilde{v},\mathcal{F})$ pointwise. Now the proof is complete. ∎
@@ B.3 Invertibility condition
Let $A$ be an SPD matrix and define the $||\cdot||_{A}$ norm $||X||_{A}:=||A^{1/2}XA^{-1/2}||_{2}$. Then for an arbitrary matrix $Y$ the inequality $||I-YA||_{A}<1$ implies that $Y$ is invertible.
@@
$||I-YA||_{A}=\left|\left|I-A^{1/2}YA^{1/2}\right|\right|_{2}=\max_{u\neq 0}\frac{\left|\left|(I-A^{1/2}YA^{1/2})u\right|\right|_{2}}{||u||_{2}}$ (B.7)
@@
$\frac{\left|\left|(I-A^{1/2}YA^{1/2})u\right|\right|_{2}}{||u||_{2}}=1,$ (B.8)
which contradicts the assumption $||I-YA||_{A}<1$. ∎

### v3

@@ A.6 Bounds on the IACT
To show the last statement, note that if $\theta_{L}^{(0)}\sim\mathcal{N}(\mu_{L},A_{L}^{-1})$ then according to Theorem 4.1 $\theta_{L}^{(m)}\sim\mathcal{N}(\mu_{L},A_{L}^{-1})$ for all $m\geq 0$. With (4.26c) in Lemma 4.10 this implies that
@@
Inserting this into (4.27) gives the desired result. To derive a bound on $\tau^{(\infty)}_{\mathrm{int},F_{L}}$, i.e. to show the first inequality in (4.29), observe that the individual terms in the sum in (4.28) can be bound as follows if we set $w:=A_{L}^{-1/2}F_{L}$
[…]
where the energy norm $||\cdot||_{A_{L}}$ is defined in (4.19). Using (A.32) and summing the geometric series we get the desired bound on $\tau^{(\infty)}_{\mathrm{int},F_{L}}$:
[…]
@@
$\tau^{(\infty)}_{\mathrm{int},F_{L}}$ $=1+2\sum_{s=1}^{\infty}\frac{\alpha_{s}}{\beta},$ $\tau_{\mathrm{int},F_{L}^{\top}}^{(m)}$ $=1+2\sum_{s=1}^{\infty}\frac{\alpha^{\prime}_{m,s}}{\beta^{\prime}_{m}}=\tau^{(\infty)}_{\mathrm{int},F_{L}}+2\sum_{s=1}^{\infty}\left(\frac{\alpha^{\prime}_{m,s}}{\beta^{\prime}_{m}}-\frac{\alpha_{s}}{\beta}\right).$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
with the grid-independent constant $C:=4(1+C_{0})$. We conclude that $\tau_{\mathrm{int},F_{L}}^{(m)}$ converges to $\tau^{(\infty)}_{\mathrm{int},F_{L}}$ for $m\rightarrow\infty$. ∎
@@ A.8 Approximation property for perturbed matrix
where we used Assumption 1.
@@
[…]
@@
Finally, Assumption 1 implies $\|P_{\ell}^{-1}\varphi_{\ell}\|_{2}\leq\frac{c_{2}}{\Phi(\ell)}\|\varphi_{\ell}\|_{H}$ for $\varphi_{\ell}\in V_{\ell}$ and thus
@@ A.9 Optimality of MGMC
Finally, noting that $u:=\widetilde{A}^{-1}f$ satisfies the problem $a(u,v)+b(u,v)=\langle f,v\rangle_{H}$ for all $v\in H$, Assumption 2 implies that $P_{L}u_{L}:=P_{L}\widetilde{A}_{L}^{-1}f_{L}$ with $f_{L}=(\langle f,\phi_{j}^{L}\rangle_{H})_{j=1,\dots,n_{L}}$ converges to $u$ in $V$ as $L\to\infty$. Hence,
@@
[…]
@@ B.3 Invertibility condition
[…]
@@
[…]
@@
[…]
which contradicts the assumption $I-YA||_{A}<1$. ∎
Appendix C Derivation of additional complexity estimates
In this appendix we present detailed derivations of the MGMC setup costs in Section 3.6.1 and storage requirements in Section 3.6.2.
C.1 Derivation of MGMC setup costs
Proof of Theorem 3.5.
In the setup phase, the Gibbs-sampler in Alg. 3 requires $\mathcal{O}(\beta n_{\ell})$ operations to solve the $\beta$ triangular systems in line 7. The computation of the dense $\beta\times\beta$ matrix $\Gamma+B_{\ell}^{\top}C_{\ell}$, its factorisation and the construction of the $n_{\ell}\times\beta$ matrix $G_{\ell}$ in line 8 incur costs of $\mathcal{O}(\beta^{2}n_{\ell})$, $\mathcal{O}(\beta^{3})$ and $\mathcal{O}(\beta^{2}n_{\ell})$ respectivly. This leads to the following bound of the setup costs for Alg. 3:
$\mathrm{Cost}^{\mathrm{(setup)}}_{\mathrm{Gibbs}}(n_{\ell},\beta)\leq C_{\mathrm{Gibbs}}^{(\mathrm{setup},1)}\beta^{3}+C_{\mathrm{Gibbs}}^{(\mathrm{setup},2)}\beta^{2}n_{\ell}.$ (C.1)
As explained in Section 3.6, when applying the low-rank correction in the multiplication by $\widetilde{A}_{\ell}$, which is required in the residual calculation, it is advantageous to precompute the $\beta\times n_{\ell}$ matrix $\Gamma^{-1}B_{\ell}^{\top}$ at a cost of $\mathcal{O}(\beta^{p_{\Gamma}}n_{\ell})$, which incurs additional costs of
$\mathrm{Cost}^{(\mathrm{setup})}_{\mathrm{other}}(\beta,n_{\ell})\leq C^{\mathrm{(setup)}}_{\mathrm{other}}\beta^{p_{\Gamma}}n_{\ell}$ (C.2)
on each level. The setup cost on the coarsest level can be bounded as
[…]
As in the proof of Theorem 3.2, we can bound the setup costs on each level $\ell$ of the hierarchy recursively
$\mathrm{Cost}^{(\mathrm{setup})}_{\mathrm{MGMC}}(\ell)\leq\begin{cases}C^{(\mathrm{setup})}_{\mathrm{other}}\beta^{p_{\Gamma}}n_{\ell}+\mathrm{Cost}^{(\mathrm{setup})}_{\mathrm{Gibbs}}(n_{\ell},\beta)+\mathrm{Cost}^{(\mathrm{setup})}_{\mathrm{MGMC}}(\ell-1)&\text{for $\ell>0$}\\ \mathrm{Cost}^{(\mathrm{setup})}_{\mathrm{coarse}}(\beta,n_{0})&\text{for $\ell=0$}\end{cases}$ (C.4)
Inserting the bounds in (C.1), (C.2) and (C.3) into this expression leads to
[…]
$\leq C_{\mathrm{MG}}^{(\mathrm{setup},1)}\beta^{3}L+C_{\mathrm{MG}}^{(\mathrm{setup},2)}\beta^{2}n_{L}+\mathrm{Cost}_{\mathrm{coarse}}^{(\mathrm{setup})}(\beta,n_{0})$
where, since $p_{\Gamma}\leq 2$ and $\rho_{G}<1$,
$C_{\mathrm{MG}}^{(\mathrm{setup},1)}$ $=C_{\mathrm{Gibbs}}^{(\mathrm{setup},1)},$ $C_{\mathrm{MG}}^{(\mathrm{setup},2)}$ $=\frac{C_{\mathrm{oher}}^{(\mathrm{setup})}+C_{\mathrm{Gibbs}}^{(\mathrm{setup},2)}}{1-\rho_{G}}.$ (C.6)
C.2 Derivation of MGMC memory requirements
Proof of Theorem 3.6.
On each level $\ell$ we need to store no more than $K$ vectors of size $n_{\ell}$, such as $\theta_{\ell}$, $f_{\ell}$, $\psi_{\ell}$ and $\xi_{\ell}^{\mathrm{diag}}$; the exact value of $K$ (which is independent of $\ell$) depends on details of the implementation. To apply Alg. 3 and to compute the residual, we need to store the sparse matrix $A_{\ell}$, which also has $\mathcal{O}(n_{\ell})$ entries. In addition, we require the storage of the vector $\xi_{\ell}^{\mathrm{LR}}$ of length $\beta$, as well as the $n_{\ell}\times\beta$ matrices $B_{\ell}$, $G_{\ell}$ and the $\beta\times n_{\ell}$ matrices $B_{\ell}^{\top}$, $\Gamma^{-1}B_{\ell}^{\top}$. This results in the following bound
$\mathrm{Memory}_{\mathrm{Gibbs}}(\ell)\leq C_{\mathrm{Gibbs}}^{(\mathrm{mem})}(1+\beta)n_{\ell}+\beta$ (C.7)
for some constant $C_{\mathrm{Gibbs}}^{(\mathrm{mem})}$. On the coarsest level, we either need to store the quantities that are required for Alg. 3 with $\ell=0$ or the Cholesky factorisation of the $n_{0}\times n_{0}$ matrix $\widetilde{A}_{0}$. This implies that
[…]
Putting everything together, the total storage cost for Alg. 1 can be bounded as follows:
$\mathrm{Memory}_{\mathrm{MGMC}}(L)$ $\leq C_{\mathrm{Gibbs}}^{(\mathrm{mem})}(1+\beta)\sum_{\ell=1}^{L}n_{\ell}+\beta L+\mathrm{Memory}_{\mathrm{coarse}}(\beta,n_{0})$ (C.9)
$\leq C_{\mathrm{Gibbs}}^{(\mathrm{mem})}(1+\beta)n_{L}\sum_{\ell=1}^{L}\rho_{G}^{L-\ell}+\beta L+\mathrm{Memory}_{\mathrm{coarse}}(\beta,n_{0})$
$\leq C_{\mathrm{MG}}^{(\mathrm{mem})}(1+\beta)n_{L}+\beta L+\mathrm{Memory}_{\mathrm{coarse}}(\beta,n_{0})$
with
$C_{\mathrm{MG}}^{(\mathrm{mem})}=\frac{C_{\mathrm{Gibbs}}^{(\mathrm{mem})}}{1-\rho_{G}}.$ (C.10)
Appendix D Additional numerical results
While in Sec. 5 we assumed that the entries of the covariance matrix $\Gamma$ lie in the interval $\widehat{\sigma},2\widehat{\sigma}$] with $\widehat{\sigma}=10^{-6}$, we also measured the integrated autocorrelation time (IACT) of the Gibbs- and MGMC sampler for a wide range of other values of $\widehat{\sigma}$ and for three different correlation lengths $\kappa^{-1}$. The results in Tab. 2 were obtained for the FEM discretisation of the shifted Laplace operator $\mathcal{A}^{(\mathrm{SL})}=-\Delta+\kappa^{2}I$ on a two-dimensional grid of size $128\times 128$. All other parameters are the same as described in Sec. 5.1. As the table shows, the IACT is approximately independent of value of $\widehat{\sigma}$. While for the Gibbs sampler it increases as the inverse correlation length grows (this is consistent with the deteriorating convergence rates in Fig. 2), for the MGMC sampler the IACT is smaller than $1.6$ for all values of $\kappa^{-1}$ and $\widehat{\sigma}$ considered here.
Gibbs MGMC
$\widehat{\sigma}$ $\kappa^{-1}=0.10$ $\kappa^{-1}=0.25$ $\kappa^{-1}=0.50$ $\kappa^{-1}=0.10$ $\kappa^{-1}=0.25$ $\kappa^{-1}=0.50$
$10^{-10}$ $41.4\pm 10.6$ $59.6\pm 17.4$ $95.8\pm 33.3$ $1.20\pm 0.13$ $1.41\pm 0.12$ $1.42\pm 0.13$
$10^{-8}$ $38.0\pm 9.4$ $88.8\pm 30.0$ $68.9\pm 21.2$ $1.23\pm 0.14$ $1.36\pm 0.11$ $1.47\pm 0.14$
$10^{-6}$ $34.6\pm 8.3$ $58.5\pm 17.0$ $118.1\pm 44.2$ $1.24\pm 0.14$ $1.42\pm 0.13$ $1.59\pm 0.12$
$10^{-4}$ $47.0\pm 12.6$ $89.1\pm 30.1$ $69.9\pm 21.6$ $1.18\pm 0.12$ $1.39\pm 0.12$ $1.55\pm 0.16$
$10^{-2}$ $35.8\pm 8.7$ $74.7\pm 23.6$ $88.0\pm 29.6$ $1.23\pm 0.14$ $1.40\pm 0.12$ $1.42\pm 0.13$
11 $46.5\pm 12.5$ $02.8\pm 36.6 11$$0.9\pm 40.6 1.1$$3\pm 0.11 1.1$$4\pm 0.11 1.1$$15\pm 0.12$
$10^{2}$ $48.2\pm 13.1$ $107.8\pm 39.1$ $50.2\pm 61.3 1.1$$3\pm 0.11 1.1$$3\pm 0.11 1.1$$15\pm 0.11$
$10^{4}$ $48.2\pm 13.1$ $107.8\pm 39.1$ $50.2\pm 61.4 1.1$$3\pm 0.11 1.1$$3\pm 0.11 1.1$$15\pm 0.11$
Table 2: IACT for different values of the correlation length $\kappa^{-1}$ and magnitude $\widehat{\sigma}$ of the covariance matrix $\Gamma$. All results were obtained with the FEM discretisation of the shifted Laplace operator $\mathcal{A}^{(\mathrm{SL})}=-\Delta+\kappa^{2}I$ on a grid of size $128\times 128$.